# Mining higher-order triadic interactions

### v3

@@ IV.2 Validation of the TRIM algorithm on the triadic interaction model
To demonstrate the efficacy of the TRIM algorithm we first considered a single network (see Figure 4$(a)$) of $N=10$ nodes, $L=12$ edges and $\hat{L}=5$ triadic interactions (each formed by a single node regulating a single edge) on top of which we consider the triadic dynamical model proposed in Sec. III.
@@
Simulations of Eq. (7) on this network show a strong dependence of $MIz(m)$ on $m$ for the triples of nodes involved in triadic interactions, with greater significance for smaller values of $\alpha$. Figure 3 shows the joint distributions $P_{\alpha}(X,Y)$ of $X$ and $Y$ for each interval $\alpha$ of values $Z$ determined by the decision tree. The results provide evidence of this interesting dynamical behavior of the triadic model in the case of a positive regulatory interaction. Note that our analysis of the function $MIz(m)$ also allows us to distinguish between positive and negative regulatory interactions. These are associated respectively by an increase in $MIz$ for larger values of $z_{m}$ or decrease of $MIz$ In the Supplementary Figures S1-S2, we display further examples of the function $MIz$ for triadic triples. We observe that as the parameter $\Gamma$ is raised, i.e. the noise increases, which is reflected in the increased variability of the $MIz$ functional behavior.
In order to provide an overall assessment of the performance of TRIM in mining triadic interactions on this network, in Figure 4$(b)$ we display the Receiver Operating Characteristic curve (ROC curve) based on the score $\Theta_{\Sigma}$, measured with respect to the first null model which we found to perform well for a wide range of parameters. As it is apparent from the ROC curve, the TRIM algorithm performs better for higher values of $\alpha$ and smaller values of $\Gamma$ (see Figure 4$(b)$). For any values of the parameters, we notice that the false positives are more likely to involve short-range triples, i.e. triples in which the regulator node $Z$ has short (small) structural network distance with the two nodes $X$ and $Y$.
Figure 5: Performance of the TRIM algorithm on a benchmark network with triadic interactions. Each data point shows measures for a triple of nodes $X$, $Y$ and $Z$, namely $\Theta_{\Sigma}$ and CMI, the proposed $\Sigma$-score and conditional mutual information between $X$ and $Y$, respectively. The colour of each point corresponds to the value of $S$, which characterises the entropy of the triple. The synthetic data comes from a network with $100$ nodes, $110$ edges, and $25$ triadic interactions. All possible triples are represented in the scatter plot. Stars are the true triadic triples.
We conclude our discussion by testing the TRIM algorithm on a much larger model network, to demonstrate its scalability. To this end, we consider a random network of $100$ nodes, $110$ edges, and $25$ triadic interactions (see SI for details). In Figure 5 we report the results of the TRIM algorithm for all the triples involving all edges, $14$ of which come from the pool of edges with triadic interactions. For each edge, all possible triples are analysed while for each edges we retain only the $5$ most significant triples according to $\Theta_{\Sigma}$. By conditioning on the value of third node, for each of these connected nodes we also record the conditional mutual information CMI. In figure 5, each considered triple corresponds to a point, colour coded according to the value of $S$. Stars indicate triples that are involved in a triadic interactions (see SI for details). The true triadic interactions are indeed found for triples with high $\Theta_{\Sigma}$ score, while CMI and $S$ span between high and intermediate values.
@@
Figure 6: Maximal Spanning Tree of the relevant genes from the gene expression data. Edges and their edge weights are obtained from the Protein-Protein interaction network. Red-coloured nodes are nodes with biological significance, that is that play a critical role in AML[47][48][49]. Node size is proportional to the node degree.
@@
Figure 7: (A) Scatter plot which displays $\Theta_{\Sigma}$ (Y-axis) and the CMI (X-axis) value between two genes, denoted $X$ and $Y$ , conditioned by a third gene, denoted $Z$. The colour corresponds to the value of our entropy measure $S$. Here only triples are depicted that rank among the 300 highest for CMI and MI and have $p$-value 0.001 or less in the randomization null model. Circles are triples resulting from the MST, whereas squares are triples involving genes with biological relevance. Crosses are triples that have $p$-value biggger than $0.0005$ with the Gaussian version null model. (B) The results of the decision tree allow us to distinguish three intervals in the value of Z in which the conditional distribution of the gene expression of $X$ and $Y$ differs the most. Plots of these conditional distributions are provided for two exemplary triples: both display a high significance score for the triadic interaction, indicating a meaningful biological association. The top row shows three biologically relevant genes, (X = HOXB3, Y = MEIS1, Z = GLIS3, $\Theta_{\Sigma}$ = 5.2, $p_{\Sigma}$ = 0.0001, $\Sigma$ = 0.38); the bottom row shows a triples identified by the MST. (C) has $p_{\Sigma}$-value 0.0001 in the random version. (X = GATA1, Y = KLF1, Z = ETV1, $\Theta_{\Sigma}$ = 6.1, $\Sigma$ = 0.44)
@@ V Detecting triadic interactions in gene-expression data
Searching for trigenic interactions in gene-expression is a problem of major interest in biology. For instance, understanding the extent to which a modulator promotes or inhibits the interplay between a transcription factor and its target gene is crucial for deciphering gene regulation mechanisms. In order to address this question with our method, we considered gene-expression data from Acute Myeloid Leukemia (AML), extracted from the Grand Gene Regulatory Network Database [50, 51].
@@
Exhaustive mining of all potential triadic interactions among each triple of nodes in the AML dataset is computationally very demanding (it would require testing of $>260M$ triplets). Instead, we here focus on edges associated with known biophysical interactions, as identified in the human Protein-Protein Interaction network (PPI)  [50]. To do so, we considered the subgraph of the human PPI network that contains all the genes/proteins included in the AML gene expression data. The number of edges in this subgraph is $M=42,511$. We first selected a set of triples involving genes known to be associated to AML with gene X and Y connected by an edge in the PPI (see SI for details). Moreover we selected additional triples according to a criteria based on the positions of the three genes in the Maximum Spanning Tree (MST) which only includes 621 edges (see Figure 6). For each edge in the MST, connecting gene X with gene Y we consider all genes Z within a distance of 4 from both the X and Y as candidate regulatory nodes, i.e. the third node in the triple (see SI for details). For each considered triple of genes we use $\Theta_{\Sigma}$ as the significance score, with $P=5$ bins, using $\mathcal{N}=5\times 10^{3}$ realizations of the randomization null model. These results have strong correlations with the results obtained considering $\Theta_{T}$ as the significance score (see Supplementary Figures S3-S4 for comparison).
@@
Figure 7 shows the result of the TRIM algorithm for a selected set of the analysed triples with $p_{\Sigma}<0.001$. Specifically, the figure displays each triple in the plane $\Theta_{\Sigma}$ and CMI with a colour code encoding the value $S$. Note that for each selected edge only the top $5$ triples ranked according to the $\Theta_{\Sigma}$ score are depicted. Squares indicate triples chosen from biologically relevant genes for AML. Crosses indicate triples deemed insignificant in comparison with the Gaussian null model, Figures 7B and 7C provide examples of conditional distributions of two exemplary triples with high entropic score $S$. Both triples show evidence of triadic interactions: the triple in panel B is a member of the MST, while C is an example of a triple chosen from known biologically relevant genes for AML. For further examples of results obtained on more gene-expression triples, see the SI for details. Interestingly, among the significant triples, we detected also triples in which the modulation of the mutual information is non-monotonic (see SI for details).
@@
Many of the genes involved in the $50$ highest ranking triples have already been linked to AML in the literature (see SI for Table 2 for a list of highly significant triples and Figure S6 with links to the literature associating the involved genes with AML). Interestingly, genes such as PAX4, SOX14 and VAX2 which appear in multiple triples have been found to be over- or under-expressed in patients with AML. Remarkably, 2 such genes, ASCL1 and NEUROG2, are involved in $5$ triples, suggesting that they may play a particularly important part in disease progression [52]. In total, $61\%$ of genes in the triples have a known association with AML.
@@ VI Conclusions
This work provides a comprehensive information theory-based framework to model and mine triadic interactions directly from dynamic observations and for previous knowledge of the underlying pairwise network. The generative model we propose demonstrates that the presence of a triadic interaction leads to the modulation of the mutual information between the two end nodes of an edge (X and Y) by a third regulatory node Z. Via this model we show that to detect triadic interactions it is necessary to go beyond standard pairwise measures, such as the mutual information. Moreover, also standard higher-order statistical measures such as the conditional mutual information that only account for the average effect of the third regulatory node Z on the mutual information between X and Y falls short in characterizing triadic interactions. Our proposed approach, implemented in the TRIM algorithm, mines triadic interactions based on the quantification of the statistically significant modulation of the mutual information between the two linked nodes conditioned on the third regulator node. This algorithm is tested and validated on our generative model using paradigmatic random networks with triadic interactions and used to mine putative triadic interactions from gene expression data.
From the network theory point of view, this work opens new perspectives in the active field of modelling and inference of higher-order interactions and can be extended in many different directions, for instance by exploring the effect of triadic interactions on the dynamical state of nodes associated with discrete variables. From the biological point of view, our results may inspire further information-theoretic approaches to genetic regulatory network inference. Investigating the extent to which triadic interactions are tissue-specific, and if certain regulatory patterns are conserved across different tissues, could yield valuable insights. Our proposed approach could also be used to mine triadic interactions in other domains, such as finance or climate, where triadic interactions also have a significant role.
@@
The Python package triaction is available on GitHub at the following link: https://github.com/anthbapt/triaction

### v4

@@ IV.2 Validation of the TRIM algorithm on the triadic perceptron model
In order to discuss the phenomenology of the TPM we first considered a representative network (see Figure 3) of $N=10$ nodes, $L=12$ edges and $\hat{L}=5$ triadic interactions (each formed by a single node regulating a single edge) on top of which we consider the TPM proposed in Sec. III.
We found that data obtained from the TPM on this network shows a strong dependence of $MIz(m_{z})$ on $m_{z}$ for the triples of nodes involved in triadic interactions, with greater significance for smaller values of $\alpha$. Figure 3 shows the difference between the $MIz(m)$ profile of a triple that is involved in a triadic interaction compared to a triple that is not, demonstrating how triadic interactions modulate the $MIz(m_{z})$ profile.
@@
Moreover, Figure 4 shows, for a given triadic interaction involving nodes X, Y and Z, the joint distributions $\mu_{\delta}(X,Y)$ of $X$ and $Y$ for each interval $\delta$ of values $Z$ determined by the decision tree. The results provide evidence of this interesting dynamical behavior of the triadic model in the case of a positive regulatory interaction. Note that the analysis of the form of the function $MIz(m_{z})$ also allows us to distinguish between positive and negative regulatory interactions, which are associated with an increase or a decrease in $MIz$ for larger values of $m_{z}$ respectively.
In the Supplementary Figures S1-S2, we display further examples of the function $MIz$ for triadic triples. We observe the increased variability of the $MIz$ functional behavior as the parameter $\Gamma$ is raised, i.e. the noise increases.
These results confirm the main general principle on which the TRIM algorithm is based, i.e. that the conditional mutual information $MIz$ is modulated by triadic interactions. To make this observation precise, we examined the performance of the TRIM algorithm in mining triadic interactions from synthetic data. We first considered the network shown in Figure 3, and using the score $\Theta_{\Sigma}$, we evaluated the Receiver Operating Characteristic (ROC) curve and Precision Recall (PR) curve for different values of the dynamical parameters (see Figure 5). Both the ROC curve and the PR curve (which addresses the limitations of the ROC curve for imbalanced datasets) indicate that the TRIM algorithm performs well on data produced by the TPM, with a better performance for higher values of $\alpha$.
For all parameter values, we noticed that false positives are more likely to involve short-range triples, i.e. triples in which the regulator node $Z$ is close (in the structural network) to the end-points $X$ and $Y$ of the target edge.
Figure 6: Performance of the TRIM algorithm on a random network with triadic interactions. (a) Each data point represents a given triple of nodes $X$, $Y$ and $Z$. The $y$-axis shows $\Theta_{\Sigma}$, while the $x$-axis shows the CMI between $X$ and $Y.$ The colour of each point corresponds to the value of $S$, which characterizes the entropic score of the triple. The synthetic data comes from a structural random Erdös-Renyi network with $100$ nodes, and average degree $c=4$, to which $25$ triadic interactions between random edges and random nodes have been added. For the modelling of the network we used $\alpha=0.06,\Gamma=1.4\times 10^{-2},t_{max}=1500,w_{+}=18,w_{-}=0.2$ and for the analysis with TRIM we looked at $3000$ data points and $P=100$ bins. We display top 5 triples for each edge according to $\Theta_{\Sigma}$ that are below our $p$ threshold for the randomization null model. The Triples below are represented in the scatter plot and they all display an entropic score $S>0.5$. Stars are the true triadic triples which are characterized by high $\Theta_{\Sigma}$. Crosses are the triples that can be excluded by performing TRIM with the Gaussian Null model. (b) Histogram of the $\Theta_{\Sigma}$-values for all the triples of the network (in light blue), and for the triples corresponding to the $25$ true triadic interactions only (in dark blue). (c) Histogram of the $\Theta_{\Sigma}$-values observed in a network of the same topology and with the same dynamical parameters for which all the triadic interactions have been removed (orange).
@@
Figure 7: Maximal Spanning Tree of the relevant genes from the gene expression data. Edges and their edge weights were obtained from the Protein-Protein interaction network. Red-coloured nodes are nodes with biological significance, that is that play a critical role in AML [47, 48, 49]. Node size is proportional to the node degree.
@@
Figure 8: Application of the TRIM algorithm to gene expression data. Panel (a) shows the results of TRIM for the significant triples in the AML dataset. The scatter plot shows $\Theta_{\Sigma}$ ($y$-axis) versus the CMI ($x$-axis). The colour of each point corresponds to the value of its entropic score $S$. Here we display only those triples with pp-value 0.001 or less in the randomization null model and that have not been excluded by the Gaussian null model (for details about these triples see SI). Circles are triples whose links all appear in the minimum spanning tree, and squares indicate triples involving genes with biological relevance. Panels (b)-(c) display the conditional distributions for two example triples: both are identified by the TRIM algorithm with high significance, suggesting a meaningful biological association. Panel (b) shows the triple $X$ = GATA1, $Y$ = KLF1, $Z$ = ETV1. According to the randomized surrogate null model, this triadic interaction has $p_{\Sigma}$-value 0.001, $\Theta_{\Sigma}$ = 4.75, $\Sigma$ = 0.44, $S=0.64$; panel (c) shows the results for the triple $X$ = HOXB3, $Y$ = MEIS1, $Z$ = GLIS3 involving two biologically relevant genes. According to the randomized surrogate null model, it has $\Theta_{\Sigma}$ = 3.98, $p_{\Sigma}$ = 0.001, $\Sigma$ = 0.38, $S=0.60$.
These results indicate that the TRIM algorithm is effective in identifying triadic interactions in a small network generated using the TPM. To examine the scalability of this methodology we also tested the TRIM algorithm on a much larger model network. To this end, we consider a random Erdös-Renyi network of $100$ nodes, and average degree $c=4$ to which we added $25$ random triadic interactions (i.e., between randomly chosen nodes to randomly chosen edges), imposing the condition that each edge is at most regulated by a single node for simplicity. The results of this analysis are shown in Figure $6(a)$, in which we provide statistics for all possible node triples in the network (the majority of which are not triadic interactions). For each edge, we retained only the $5$ most significant triples according to $\Theta_{\Sigma}$. By conditioning on the value of third node, for each of these connected nodes we also record the conditional mutual information CMI. In Figure 6, each considered triple corresponds to a point, colour coded according to the value of $S$. Stars indicate triples that are involved in a triadic interactions (see SI for details). True triadic interactions are found for triples with high $\Theta_{\Sigma}$ while CMI span between high and intermediate values. This result confirms the very good performance of the TRIM algorithm on the data coming form the TPM.
To test the statistical robustness of the TRIM algorithm we also conducted the same analysis (i.e., on the same structural network with the same dynamical parameters) in which all triadic interactions were removed. The results of this analysis are shown in Figure 6(c). In this case, and as expected, the TRIM algorithm did not identify any statistically significant triadic interactions. This analysis indicates that the TRIM algorithm is able to identify true triadic interactions with a low false positive discovery rate (compare Figures 6(b)-(c)).
@@ V Detecting triadic interactions in gene-expression data
Searching for triadic interactions in gene-expression is a problem of major interest in biology. For instance, understanding the extent to which a modulator promotes or inhibits the interplay between a transcription factor and its target gene is crucial for deciphering gene regulation mechanisms [29]. In order to address this question with our method, we considered a gene-expression dataset associated with Acute Myeloid Leukemia (AML), extracted from the Grand Gene Regulatory Network Database [50, 51].
Exhaustive mining of all potential triadic interactions from every putative triple of nodes in the AML dataset is computationally very demanding (it would require testing of >260M>260M triples) and likely, due to the sheer number of triples being tested, to result in false positives and/or interactions of less biological importance. Moreover, such a brute-force approach would not account for other sources of important biological information, such as putative interactions derived from other experimental sources. To account for such information, we therefore focused our analysis on edges between nodes associated with known biophysical interactions, as identified in the human Protein-Protein Interaction network (PPI) [50]. To do this, we considered the connected subgraph of the human PPI network that contains all the genes/proteins included in the AML gene expression data and their associated edges This network, which contains $622$ nodes and $42,511$ edges, formed the structural network for our analysis [52]. To start, we focused on triples involving genes known to be associated with AML, in which the end-points $X$ and $Y$ of the target are directly connected in the PPI network (see SI for details).
@@
We then selected additional triples according to their positions in the PPI network’s Maximum Spanning Tree (MST), which only includes 621 edges (see Figure 7). In order to focus on triples for which network effects are likely to be less pronounced, for each edge in the MST, connecting gene/protein $X$ with gene/protein $Y$ we considered all genes $Z$ within a distance of 4 from both the $X$ and $Y$ as candidate regulatory nodes, i.e. the third node in the triple (see SI for details). For each considered triple of genes we assessed its significance using $\Theta_{\Sigma}$ as the significance score, with $P=5$ bins, using $\mathcal{N}=5\times 10^{3}$ realizations of the randomization null model (very similar results were obtained using $\Theta_{T}$ as the significance score, see Supplementary Figures S3-S4 for a comparison).
@@
Figure 8 shows the results of the TRIM algorithm for those triples with $p_{\Sigma}<0.001$. Note that for each selected edge only the top $5$ triples ranked according to the $\Theta_{\Sigma}$ score are depicted. Squares indicate triples chosen from biologically relevant genes for AML. The triples deemed insignificant according to the Gaussian null model, are not shown here. The interested reader can see their visualization in Figure $S6$ of the SI. Figure 8 provide examples of conditional distributions of two illustrative triples that rank high in TRIM. Both triples show evidence of a triadic interaction: the triple in panel (b) is a member of the MST, while (c) is an example of a triple chosen from known biologically relevant genes for AML. Further example triples are shown in the SI. Interestingly, among the significant triples, we detected also triples in which the modulation of the mutual information is non-monotonic (see SI for details).
@@
Many of the genes involved in the $50$ highest ranking triples have already been linked to AML in the literature (see SI for Table S3 for a list of highly significant triples and Table S4 with links to the literature associating the involved genes with AML). In total, $84\%$ of the top 50 Triples include at least one gene that has a known association with AML.
@@ VI Conclusions
This work provides a comprehensive information theory-based framework to model and mine triadic interactions directly from dynamic observations. The TPM we propose demonstrates that the presence of a triadic interaction leads to systematic variations in the mutual information between the two end nodes of the edge involved ($X$ and $Y$). Via this model we have shown that to detect triadic interactions it is necessary to go beyond standard pairwise measures, such as the mutual information. Importantly, standard higher-order statistical measures, such as the conditional mutual information, which accounts for the average effect of the third regulatory node $Z$ on the mutual information between the target nodes $X$ and $Y$ are also insufficient to identify triadic interactions. Our proposed approach, implemented in the TRIM algorithm, mines triadic interactions by identifying statistically significant variations in the mutual information between the two linked nodes conditioned on the third regulator node.
To demonstrate the efficacy of this algorithm we have tested and validated it on a new dynamical model (that we denote the TPM) and shown how it can identify triadic interactions in randomly generated triadic interaction networks. We also used it to mine putative triadic interactions from gene expression data, and connect the putative interactions with meaningful biology.
From the network theory point of view, this work opens new perspectives in the active field of modelling and inference of higher-order interactions and can be extended in many different directions, for instance by exploring the effect of triadic interactions on the dynamical state of nodes associated with discrete variables or including time delays in the regulation. From the biological point of view, our results may inspire further information-theoretic approaches to genetic regulatory network inference. Investigating the extent to which triadic interactions are tissue-specific, and if certain regulatory patterns are conserved across different tissues, could yield valuable insights. Our proposed approach could also be used to mine triadic interactions in other domains, such as finance or climate, where triadic interactions also have a significant role.
@@
The Python package TRIM is available on GitHub at the following link: https://github.com/anthbapt/TRIM